Shift Summary for Day shift, 06/03/99

Jianbo Wang (jianbo@fnal.gov)
Thu, 3 Jun 1999 15:39:43 -0500 (CDT)

Shift Summary for Day shift, 06/03/99

On Shift: Hogan
Jianbo
A variety of other committed collaborators were working throughout this
shift.

Hours of beam delivered at all intensities: 6.5
Hours of beam delivered at nominal intensity: 6.0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): hrs

Accounting of lost time:

There is warning message said that home disk of ktev3b is 98% full and the
limit of it is 95%. Hogan fix it by moving some log file to data disk.
The culprits are large log files generated by ktevana errors occuring
during the run in question. See below.

10:50am We got a lot of ktevana errors at ktev1 (which generate large disk
files), there were the same error last night, but the ktevana monitor
histo didn't show those errors. Jason L fixed by replacing a faulty
memory of plane 1.

11:00 We got 4225 ktevana errors of missing FFFF in stream 5. Stop the run
and start the new run, the problem's gone. It's caused crate 53.

11:18 RT, AT, Leo : Make control access to remove mask, change the
threshold of HA1 to 261mv,
threshold of HA2 to 550mv.

11:22 SET-RC9 src is back. Tight CA-L handle.

BobT called control room to change magnet to E799 configuration. Arun
JasonH, JasonL started 799 run to study trigger.

15:00 Make control access to push last TRD out of beam.

Please call Glen at 4836 for beam stop monitor tomorrow after the beam is
down.

========================================================================
Runs ended this shift:
Run 13324 - completed
Run 13325
Run 13326 -

Tapes brought to FCC this shift: 0